### v1

# Machine Learning to Predict Spectral Anisotropy in Valence-to-Core X-ray Emission Spectroscopy

Machine Learning to Predict Spectral Anisotropy in Valence-to-Core X-ray Emission Spectroscopy
Charles A. Cardot, John Tichenor, Seth M. Shjandemaar,
Josh J. Kas, Fernando D. Vila, Gerald T. Seidler, John J. Rehr
(*) jjr@uw.edu
Polarization analysis in x-ray spectroscopy provides an orientation dependent sensitivity to local bonding environments. For a cluster of atoms, polarization sensitivity is most often discussed through the lens of point group symmetries. However, this is a discrete, qualitative method of classifying clusters, and it does little to indicate the degree of spectral anisotropy. Here we adopt a random forest model for quantitative prediction of spectral anisotropy. Its input relies on simplified local geometric and chemical information that can be obtained from any crystal structure file. The model is trained on over 10,000 experimentally realized transition metal structures from the Materials Project, with the target being VtC-XES calculated using the real space Green’s function code FEFF. We find that the model can strongly predict the degree of spectral anisotropy, with the primary factors being derived from the spatial moments of ligands in a cluster.
1 Introduction
Anisotropic materials have a multitude of applications in a wide variety of fields, including aerospace engineering, high speed electronics, thermo-, photo-, and piezo-electronics, and more [chen_2022, gong_2017, li_2021]. These materials exhibit directional dependence in mechanical, electronic and magnetic properties, making them highly desirable for use in a wide variety of devices, including optical components and polarization detectors, high-speed transistors, lightweight load-bearing devices, and energy storage solutions to name a few. There are many techniques for measuring the anisotropy of a given property [loudon-1964, cahill-2003, ziman-1972]. However, x-ray spectroscopy and related techniques are unique in their ability to probe the anisotropy of the local atomic, magnetic, and electronic structure of the material, allowing analysis of the fundamental structure-function relationships [tortora-2024].
Polarization is prevalent in x-ray spectroscopy as a method of analyzing spin and orbital magnetic moments [van-der-laan-2014], magnetic axes or interfaces [jo-2004, van-der-laan-2011], crystal field effects [bergmann-2002], and coordination environments [jansing-2016]. With the advent and improvement of high brilliance synchrotron radiation over the last 40 years, a diverse suite of experimental capabilities have emerged that can use polarization as an additional probe within x-ray spectroscopy to extract electronic and chemical information [glatzel-2004, pollock-2015, ketenoglu-2022, trippe-2014, stohr-1992]. In x-ray absorption spectroscopy (XAS) it can be used as a spin-dependent probe of the unoccupied density of states via x-ray magnetic circular dichroism (XMCD) [stohr-1999], or as a probe of bond directions and magnetic axes in transition metal oxides through x-ray linear dichroism (XLD) [bianconi-1988].
[…]
Figure 1: (a, c) Stretched NiO cluster, with the Ni (silver) to O (red) bond along the z-axis stretched 5% and 50% respectively. (b, d) The Ni VtC-XES polarized along the xy bonds in green and z bond in orange for the clusters in (a) and (c) respectively.
Local anisotropy is regularly treated in terms of discrete classifications, namely point group symmetries [watanabe-1966, brener-2024, nielsen-2024]. These are used in conjunction with oxidation state to interpret the ground state electron configuration for transition metal oxides in a given crystal field (ex: high-spin and low-spin) [lafuerza-2020]. However, it is difficult to predict a priori how large of an effect the point group derived crystal field will have on a general x-ray spectrum. This is illustrated in Figure 1, which shows how the polarization dependent VtC-XES changes when going from a 5% stretch of the apical O bond-length (a, b) to a 50% stretch (c, d). While the extent of broken octahedral (Oh) symmetry is vastly different, these clusters have the same point group (D4h).
Thus, there is a need for a continuous, quantitative metric that encodes local anisotropic environment in a way that can be used for predicting polarization dependence. Such a metric would enable rapid interpretation of experimental data and support the development of artificial intelligence (AI) driven workflows at synchrotron facilities. AI-based tools have already been successfully applied to phase identification, as well as parameter extraction in x-ray diffraction (XRD) and x-ray reflectometry (XRR), enabling continuous and rapid data-to-decision pipelines for beamline science [pithan-2023]. This metric can also be used for bulk, quantitative screening of large materials databases to identify candidates with strong local anisotropy effects. This kind of targeted exploration speeds up the search for functional materials in areas like spintronics [zutic-2004, wang-2023, tombros-2008], quantum information [shen-2022, hirohata-2025], and energy conversion and storage [sendek-2018].
Recent advances in ab initio theoretical methods have made it possible to reliably calculate a wide range of x-ray spectroscopies [vinson-2011, apra-2020]. Despite these successes, such calculations remain computationally demanding and are often limited by system size and the need to treat electron correlation [pople-1999]. By contrast, a trained machine learning model can make predictions orders of magnitude faster, enabling rapid exploration of material spaces.
The relationship between a material’s local chemical and geometric environment, and its polarization-dependent spectral response, is inherently nonlinear and highly complex. Spectral anisotropy arises from directional variations in the electronic states, which are influenced by factors such as local bonding asymmetry, ligand field distortions, coordination number, and electronic hybridization [Drager.Brummer1984]. Traditional linear or descriptor-based approaches struggle to capture these coupled, high-dimensional effects. Machine learning models, by contrast, are well-suited to learn nonlinear mappings from structural and chemical features to spectral response.
Machine learning has successfully been applied to several modalities of spectroscopy including x-ray absorption near-edge structure (XANESNet) [rankine-2020], extended x-ray absorption fine structure [timoshenko-2018], and vibrational spectroscopy [gastegger-2017]. The broader field of materials sciences has also demonstrated that machine learning models can successfully predict bulk material properties, such as utilizing a message-passing graph neural network (MPGNN) to capture anisotropic properties such as elastic [wen-2024] and dielectric tensors [lou-2024]. However, these GNN approaches can be limited by the scarcity of high-quality training data and, like most neural network-based methods, suffer from a black-box nature that limits interpretability.
In contrast, random forest models have demonstrated strong performance in materials spectroscopy; for example, Zheng et al. successfully applied them to identify coordination environments in XANES spectra [Zheng.etal2020]. Random forests can capture nonlinear and interacting effects among structural and chemical features, making them well-suited for predicting complex spectral responses while maintaining interpretability. To the authors’ knowledge, anisotropy in the VtC-XES regime remains largely unexplored and has not yet been the subject of any machine learning applications. In this paper, we develop a random forest regression (RFR) model trained on structural and chemical descriptors of 3dd transition metal complexes to predict the spectral anisotropy in VtC-XES. To quantify the degree of anisotropy, we introduce a new metric termed the spectral anisotropy matrix sum (SAMS).
The rest of this article is organized as follows. In section 2 we described the procedure for quantifying the spectral anisotropy and how we extract input features as our predictors in the model. We also address the calculation pipeline used for developing the training data set, and the procedure used to develop the random forest model. In section 3 we present the core results of the model, namely the $R^{2}$ and mean absolute error (MAE). We also assess the relative importance of the features and compare predictive capabilities across the 10 3dd transition metals. Finally in section 4 we discuss constrains of the model and propose future directions for how it could be improved and extended to new applications.
2 Methods
2.1 Spectral Anisotropy Matrix
To quantify the degree of anisotropy in a spectrum we introduce a new metric based on the $L^{2}$-norm or Euclidean norm of the difference between individual polarizations [bishop-2007]. We start by breaking a spectrum into the emission for three orthogonal polarization axes ($\sigma_{i},\sigma_{j},\sigma_{k}$). Then by calculating the square root of the integral over the difference squared we arrive at a single number to characterize the ”difference” between two curves. This value is normalized by the integral of the average (isotropic) spectra, $\bar{\sigma}$. We calculate this parameter for each possible pair of axes to produce the spectral anisotropy matrix (SAM) as defined in Eq. 1.
$SAM_{ij}$ $=\frac{\left[\int_{a}^{b}|\sigma_{i}(\epsilon)-\sigma_{j}(\epsilon)|^{2}\,d\epsilon\right]^{1/2}}{\int_{a}^{b}\bar{\sigma}(\epsilon)d\epsilon}$ (1)
$SAM_{sum}$ $=SAM_{ij}+SAM_{jk}+SAM_{ik}$ (2)
The SAM is a symmetric matrix with zeroes along the diagonal. A visualization of the calculation process is shown in Figure 2, where the Cr VtC-XES is calculated along the $x$, $y$, and $z$ crystal axes of an example transition metal crystal system, LiCrCO3F2. The crystal structure is shown in (a), where the local structure around the Cr atoms is distorted such that there are no degenerate axes. The polarized spectra along the three crystal axes is shown in (b) along with the difference curves between each pair of spectra. These are then used to calculate the Euclidean norm and build the SAM shown in (c). Taking the sum of the upper off diagonal components as shown in Eq. 2, we arrive a quantitative measure, the spectral anisotropy matrix sum (SAMS). This metric captures the total amount of anisotropy in the spectrum, and is the numerical value we aim to predict with our model in this work.
A SAM is highly dependent on the basis used to define the cluster around the transition metal ion. Our goal is to create a dataset of the maximum possible SAMS signal that can be achieved for a given structure, which often requires rotating to the principal axes. To address this, we choose to operate in the basis that diagonalized the quadrupole matrix (both for calculating the SAMS and all other features used in the model), which provides a more natural basis for maximizing the spectral anisotropy. The relationship between the spectra and the quadrupole moment will be discussed in more detail in section 2.3.
Figure 2: (a) Crystal structure of LiCrCO3F2, where Cr is bonded to five O atoms and one F atom in a distorted Oh cluster. (b) Polarized VtC Cr XES along the $x$ (blue), $y$ (green), and $z$ (red) axes, with differences between each pair shown below. (c) Spectral Anisotropy Matrix (SAM), where diagonal values are zero and off-diagonal values are the Euclidean norm of the differences (|xy||x-y|, |xz||x-z|, |yz||y-z|).
2.2 Feature Anisotropy Matrices
To parameterize the anisotropy of the local atomic environment around an absorber atom, we construct 3 related but distinct metrics derived from the quadrupole moment (Eq. 3a), dipole moment (Eq. 3b), and inertia tensor (Eq. 3c) respectively.
$Q^{\mathrm{weighted}}_{ij}$ $=\sum_{n}q_{n}\,\frac{r_{n,i}\,r_{n,j}}{\lVert\vec{r}_{n}\rVert^{7}}$ (3a)
$D^{\mathrm{weighted}}_{i}$ $=\sum_{n}\frac{q_{n}\,r_{n,i}}{\lVert\vec{r}_{n}\rVert^{7}}$ (3b)
$I^{\mathrm{weighted}}_{ij}$ $=\sum_{n}\frac{r_{n,i}\,r_{n,j}}{\lVert\vec{r}_{n}\rVert^{7}}$ (3c)
Here, $q_{n}$ is the effective charge assigned to the $n$-th neighboring atom (estimated via oxidation state), $r_{n,i}$ and $r_{n,j}$ are the $i$-th and $j$-th Cartesian components of the position vector $\vec{r}_{n}$ relative to the absorber atom, and $\lVert\vec{r}_{n}\rVert$ is the Euclidean distance from the absorber to the $n$-th atom. To emphasize the influence of nearby atoms, we introduce a distance-dependent weighting factor of $1/\|\vec{r}_{n}\|^{7}$, ensuring that contributions from closer atoms dominate while those from more distant atoms decay rapidly. This steep falloff reflects the localized nature of anisotropic interactions relevant to x-ray emission and allows the resulting descriptor to better capture short-range geometric features. The choice of exponent was determined through systematic optimization over integer values ranging from 1 to 14.
The quadrupole and inertia matrices are identical with the exception of the effective charge $q_{n}$, which serves to encode chemical information. We will see that the anisotropy metrics of both are strongly correlated with the spectral anisotropy, but we choose to use the eigenvectors of the weighted quadrupole moment to define a rotation-invariant basis. This is done both to maximize the observed spectral anisotropy in Eq. 2 and provide a set coordinate system to work in. The main reason for including both the quadrupole Eq. 3a and the inertia Eq. 3c comes from the reliance on estimating the ligand charges $q_{n}$. We rely on simple oxidation state estimations and avoid doing any direct electronic structure calculations for generating input features, and uncertainties in the charge estimates can degrade the reliability of the quadrupole metric. Incorporating a purely geometric metric thus provides a complementary, charge-independent measure of local anisotropy that can help disentangle geometric effects from uncertainties in ligand charge assignment.
The quadrupole, dipole, and inertia anisotropy matrices (denotes QAM, DAM, and IAM respectively) are shown in Eq. 4. They are all expressed in the eigen-basis of the quadrupole matrix (eigenvalues $\lambda$), where $\tilde{D}=V^{-1}DV$ and $\tilde{I}=V^{-1}IV$ are the dipole and inertia matrices transformed into this basis and $V$ is the matrix whose columns are the orthonormal eigenvectors of the weighted quadrupole moment.
$QAM_{ij}$ $=|\lambda_{i}-\lambda_{j}|$ (4a)
$DAM_{ij}$ $=\lvert\tilde{D}_{i}-\tilde{D}_{j}\rvert$ (4b)
$IAM_{ij}$ $=\lvert\tilde{I}_{i}-\tilde{I}_{j}\rvert$ (4c)
Consistent with the single anisotropy measure used for summarizing the spectral anisotropy metric into the SAMS in Eq. 2, a similar sum is applied to the off diagonal elements of Eq. 4 to arrive at the quadrupole, dipole, and inertia anisotropy matrix sums, denoted QAMS, DAMS, and IAMS respectively (Eq. 5). If a unit cell contains multiple unique sites for a 3dd transition metal, we average the anisotropy matrix sum over all sites to obtain a single representative number for the entire crystal.
$QAMS$ $=QAM_{ij}+QAM_{jk}+QAM_{ik}$ (5a)
$DAMS$ $=DAM_{ij}+DAM_{jk}+DAM_{ik}$ (5b)
$IAMS$ $=IAM_{ij}+IAM_{jk}+IAM_{ik}$ (5c)
2.3 Connection between Quadrupole and Spectra
[…]
The continuous form of the quadrupole moment tensor is defined in Eq. 7 (without additional distance weighting). This gives the quadrupole moment of a charge distribution with electron density $\rho_{e}(\vec{r})$ (Eq. 8), where $e$ is the electron charge, $N$ is the number of electrons, and $\Psi$ is the many-body wavefunction.
$Q^{continuous}_{i,j}=-\int e\rho_{e}(\vec{r})r_{i}r_{j}d^{3}\vec{r}$ (7)
$\rho_{e}(\vec{r})=N\int|\Psi(\vec{r},\vec{r}_{2},...,\vec{r}_{N})|^{2}d^{3}\vec{r}_{2}...d^{3}\vec{r}_{N}$ (8)
In Eq. 6, the initial and final state wavefunctions are shaped by the local chemical environment. This manifests as polarization-dependent effects in the spectrum. Since the electron density in 8 is derived from the modulus squared of the many-body wavefunction, it retains similar spatial anisotropy, albeit without phase information. The quadrupole moment is therefore a natural option for encoding this anisotropy, capturing both geometric and electronic symmetry-breaking features of the local environment.
To further assess the relationship between the QAMS and the SAMS, we now examine how they respond to systematic geometric distortions in representative materials, specifically NiO and Cr2O3, as shown in Figure 3. Subplots (b) and (d) show the crystal structure for NiO and Cr2O3 respectively, while subplots (a) and (c) show the QAMS vs. SAMS for different degrees of stretch or compression from the experimental structure. In (a) the NiO structure is stretched along the $c$-axis in increments of 0.5%. Starting from the lower left-hand corner of the plot—where all bond lengths are equal—the structure is isotropic, and both QAMS and SAMS are exactly zero, which is in agreement with the perfect cubic NiO structure. As the structure becomes increasingly stretched, both QAMS and SAMS vary linearly. This suggests that for simple distortions such as bond elongation without significant angular rearrangement, QAMS tracks spectral anisotropy with high fidelity.
Figure 3: QAMS versus SAMS parameters for distorted crystal structures, NiO (a, b) and Cr2O3 (c, d). NiO exhibits a linear relationship between the two parameters in response to stretching, but Cr2O3 has a nonlinear response the in the region of small distortions around the base experimental structure.
However, in the case of Cr2O3 (c) we observe a distinctly nonlinear relationship between QAMS and SAMS. Here the local environment around the Cr ion is a distorted Oh, and has non-zero QAMS and SAMS for even the starting crystal structure. In the extremes of both compression and stretching, the relationship between the two parameters remains linear. In the region around the unchanged structure, we note that the relationship deviates from linearity, and that the minimum of the SAMS does not line up with the minimum of the QAMS.
This behavior leads to two key insights. First, the nonlinear behavior arises from changes not only in bond lengths but also in bond angles, which substantially affect the spatial charge distribution and, consequently, the spectral anisotropy. For NiO, the stretching occurs uniformly without altering angular geometry, but Cr2O3 has a more complex structural responses to distortion. While the two parameters are clearly correlated, their relationship is not strictly proportional. Second, it’s possible to have a local cluster that produces a zero QAMS and a non-zero SAMS (for example, a perfectly linear molecule where the ligands have the same charge q). To address this issue, we require additional features that can more effectively capture and discriminate between the sources of anisotropy. The first of these is the DAMS from Eq. 5b, which provides orthogonal information compared to the QAMS due one being derived from a metric that is odd under inversion while the other is even. A demonstration of how DAMS can help address edge cases is provided in Supplemental Information section 6.1.1
2.4 Additional Input Features
This section focuses on defining each additional descriptor beyond the anisotropy-matrix–derived quantities and motivating its inclusion in the model. As we have shown, the QAMS alone is not a strong predictor of the SAMS, which motivates incorporating additional local geometric and chemical factors. Along this line, we include the density, the normed space group number, the number of 3dd-electrons, standard deviation and average electronegativity of the cluster as additional factors. The correlation between all factors, along with the SAMS, is displayed in Figure 4.
Figure 4: Correlation matrix of the features. Each cell shows the Pearson correlation coefficient between pairs of features, with values ranging from 1-1 (perfect anticorrelation, blue) to +1+1 (perfect correlation, red).
The normalized space group number ($N_{SG}$), as shown in Eq. 9,
$N_{\mathrm{SG}}=1-\frac{N_{s}}{N_{t}}$ (9)
where $N_{s}$ is the space group number of the structure, and $N_{t}=230$ is the total number of crystallographic space groups. Eq. 9 provides a rough, simple analogue for the symmetry of the crystal system and, to a lesser extent, the symmetry of the local environment around the metal.
The number of $3d$ electrons on the central atom is given in Eq. 10,
$N_{3d}=Z_{c}-18-q_{c},$ (10)
where $Z_{c}$ is the atomic number of the central atom and $q_{c}$ is its charge, as determined by the oxidation state guesser. The constant $18$ accounts for the electrons in the filled inner shells ($1s^{2}\,2s^{2}\,2p^{6}\,3s^{2}\,3p^{6}$), and this parameter is intended to add additional chemical information in a way that is analogous to oxidation state.
The average and standard deviation of electronegativity were added as features to capture the distribution of atomic bonding tendencies within each cluster. These quantities were computed using the Pauling electronegativity values provided by the Pymatgen periodic table [Ong.etal2013]. For each cluster, we evaluated the electronegativity of all constituent atoms and then calculated the mean and standard deviation, which together characterize both the central tendency and variability of electronegativity in the system. The density of the structure is obtained from the Materials Project [Jain2013].
Overall, Figure 4 indicates that while none of the additional descriptors individually shows a dominant linear correlation with the SAMS, several exhibit weak-to-moderate correlations that are distinct (uncorrelated) from those of the matrix-based anisotropy quantities.
2.5 Dataset curation
Crystallographic Information Files (CIFs) were queried from the Materials Project in their primitive cell format [Jain2013]. We selected structures that are experimentally observed and that contain at least one 3dd transition metal atom, resulting in a dataset of approximately 11,500 compounds. Figure 6 (a) shows the distribution of transition metals within the dataset, illustrating that certain elements, such as Cu and Fe, are overrepresented relative to others like Sc or Ti, reflecting trends in the Materials Project database. Figures 6 (b) and (c) further show the distributions of the QAMS and SAMS metrics respectively. While not identical for every element, the QAMS and SAMS distributions for the same transition metal are qualitatively similar. The most spectrally isotropic element is the Sc, with very few compounds at all that exhibit any significant anisotropy. As we will see in Section 3, the limitations of the Sc subset (both the limited number of compounds and the narrow distribution) will make it difficult to for the model to sufficiently generalize. When splitting the dataset into training and testing subsets, this elemental imbalance is mitigated by selecting a proportionally representative number of structures for each transition metal.
2.6 Corvus Workflow and FEFF
Corvus is an automation workflow handler [Kas.etal2021a] which uses a CIF and user defined information to write a FEFF input file. FEFF is a real-space Green’s function code that employs a full multiple-scattering formalism [FEFF9.6] for calculating the polarized VtC-XES. The following procedure is laid out in a diagram in Figure 5. Corvus starts by using pymatgen [Ong.etal2013] to predict the oxidation states of all the atoms in a unit cell, which are in turn used to construct the weighted quadrupole matrix as defined in Eq. 3a. An important note is that for crystal structures that contain multiple instances of the same 3dd transition metal, the quadrupole tensors are averaged over all sites before diagonalizing. The orthogonal basis that comes from diagonalizing the quadrupole is then used as the basis vectors for the subsequent FEFF VtC-XES calculation, and the generation of all subsequent features.
FEFF input files are programmatically generated with individual clusters around each transition metal element in the unit cell. The SCF and FMS radii are dynamically set to contain the closest 30 atoms to the central atom in each calculation, which helps achieve consistent convergence. All calculations were performed in the presence of a core hole, consistent with previous theoretical treatments of VtC-XES and the final-state rule [mortensen-2017]. It should be noted that we also limit the number of unique potentials so that the absorbing atom in the center of the cluster is assigned one unique potential, while all other atoms have a potential which is only unique to their atomic number. The polarized output spectra are used to calculate the SAMS according to Eq. 1 and 2 to produce a target property that our model tries to predict. The total computation time across the full dataset took approximately 3300 hours of CPU time.
Figure 5: Overview diagram of the Corvus workflow used to construct the dataset. Materials were obtained from the Materials Project and restricted to compounds containing at least one 3dd transition-metal ion that are flagged as experimentally observed.
Download CIFs from Materials Project Estimate oxidation states of atoms using Pymatgen Build out clusters, diagonalize quadrupole tensor. Formulate the QAMS, use the eigenvectors as an orthogonal basis for the spectral calculation. FEFF Polarized VtC-XES Calculate integrated spectral differences according to Eq.1 and Eq.2. Compile the SAMS together into a complete dataset
Figure 6: (a) Distribution of the 3dd transition metals in the dataset, showing the total number of structures containing each metal. (b, c) Distribution of (SAMS, QAMS) grouped by transition metal.
2.7 Machine Learning
In this work, we employ a random forest regressor (RFR) [breiman2001random] to predict the SAMS of 3dd transition metal clusters. An RFR is an ensemble learning method in which many decision trees are trained on bootstrapped samples of the data [glassner-2021], and the final prediction is obtained as the average of their outputs. Each tree is built through recursive binary splitting of the feature space, where only a randomly selected subset of features is considered at each split, thereby reducing correlation among trees and improving model robustness [james2023isl, hastie2009elements, loh2014fifty]. This model architecture provides a practical compromise between accuracy, interpretability, and computational cost, making it well suited for relating the constructed features to the anisotropy observed in VtC-XES. The RFR implementation was taken from scikit-learn [sklearn-rfr].
2.7.1 Model Training and Hyperparameter Optimization
The machine learning pipeline begins with the dataset being partitioned into training and testing subsets, with $80\%$ of the data used for training and the remaining $20\%$ reserved for testing. Hyperparameters were determined using a randomized grid search with $k$-fold cross-validation. Key RFR hyperparameters include the number and depth of trees, minimum samples per split/leaf, and the number of features considered at each split. We use bootstrap sampling (with optional subsampling) and squared error as the splitting criterion. In defining the hyperparameter space $\mathcal{G}$, we constrain the RFR hyperparameter space (e.g., depth, number of trees, and minimum samples per split/leaf) to mitigate overfitting and avoid memorization of training data. The constraints are defined in Table 1.
Table 1: Hyperparameter search constraints used during optimization of the RFR model.
Hyperparameter Constraint (Search Space)
bootstrap [True]
criterion Fixed: squared_error
max_depth {6, 8, 10, 12}
max_features {"sqrt", "log2", 0.3}
max_samples Uniform(0.6, 0.8)
min_samples_leaf randint(5, 25)
min_samples_split randint(10, 50)
n_estimators randint(300, 700)
We also employ $k$-fold cross-validation to further mitigate overfitting. The final performance estimate is obtained by averaging the results across all $k$ folds. By repeatedly shuffling and partitioning the data in this manner, we obtain a more reliable estimate of the model’s generalization performance on unseen data. For 100 randomly selected hyperparameter configuration from $\mathcal{G}$, performance was evaluated using $k$-fold cross-validation with the coefficient of determination ($R^{2}$) as the scoring metric. The highest performing hyperparameters were chosen for our model and are listed in Table 2.
Table 2: Fixed hyperparameters used for the RFR model.
Hyperparameter Value
bootstrap True
criterion squared_error
max_depth 12
max_features 0.3
max_samples 0.6676
min_samples_leaf 6
min_samples_split 11
n_estimators 587
3 Results & Discussion
3.0.1 Model Results
We use the coefficient of determination $R^{2}$ [chicco2021rsq] to evaluate model performance along with the mean absolute error (MAE). The $R^{2}$ metric quantifies the fraction of variance in the target variable explained by the model predictions, while the MAE provides an alternative performance metric through the average absolute prediction error.
Table 3: Performance metrics ($R^{2}$ and MAE) for the RFR model on the training and testing sets, with $\Delta$ (test – train).
Metric Training Testing $\Delta$
$R^{2}$ 0.8505 0.8084 0.0421-0.0421
MAE 0.0260 0.0280 +0.0020+0.0020
Figure 7: Predicted vs. True SAMS values from the trained RFR model acting on the testing set. The dashed diagonal line represents the ideal case where predictions exactly match the actual values, and the color scale indicates point density.
Our model achieves a high predictive accuracy, and the small difference between training and testing performance, $\Delta R^{2}=-0.0421$ and $\Delta\mathrm{MAE}=+0.0020$, indicates that the model generalizes well to unseen data. The relatively high $R^{2}$ score ($0.81$ on the testing set) demonstrates that the model captures the majority of the variance in the SAMS from the chosen chemical and geometric input features. This level of predictive performance is consistent with other recent applications of machine learning in materials spectroscopy and property prediction, where $R^{2}$ values in the range of 0.70.70.90.9 are typically reported for regression tasks involving complex, correlated physical quantities (e.g. chemical interpretation of near-edge x-ray absorption structure) [torrisi-2020, iwayama-2022, guda-2021].
Figure 7 shows the results of the trained model run on the test dataset. The vertical axis is the predicted SAMS value while the horizontal axis is the true SAMS value. We see that the distribution is strongly skewed along the diagonal, which is consistent with the high $R^{2}$ performance. Most of the materials are weakly anisotropic or not anisotropic at all, leading to a large density of points near the origin of the plot. Despite its strong overall performance, some systematic issues are evident. The model tends to slightly under predict SAMS values at the higher end of the distribution, while predictions for smaller SAMS values exhibit higher density and tighter clustering. This bias likely reflects the imbalance in the available data (Figure 6), where extreme high-anisotropy configurations are less common.
3.0.2 Model Interpretability
RFR feature contributions can be interpreted using variable importance measures [strobl-2007], and in this work we focus on evaluation via permutation feature importance (PFI). We chose this measure because impurity-based feature importance can be biased, especially for features with many distinct values [strobl-2007]. This is relevant for our data, where SAMS values can take any real number between 0 and 1. Impurity-based measures also depend on training-set statistics and therefore may not reflect predictive performance on unseen data.
PFI measures the importance of a feature by quantifying the change in model performance after the feature has been randomly permuted. In simple terms, this means shuffling the feature’s values among all data points to break its relationship with the target and observing how much the model worsens. If performance degrades substantially, the feature is considered more important. A more detailed discussion of PFI can be found in [molnar2025].
Figure 8: Noise-normalized permutation feature importance for the Random Forest Regressor. Bars represent the mean decrease in performance (increase in MSE) when each feature is permuted, normalized by the importance of a random noise baseline (set to 1). Values greater than one indicate predictive signal beyond random noise.
A feature ($j$) is considered important if the average importance, $\overline{\text{PFI}}_{j}$, is large, as this indicates a substantial degradation in performance when the feature is permuted [molnar2025interpretable]. However, the notion of “large” depends on the scale of the evaluation metric. Because the raw magnitudes of $\overline{\text{PFI}}_{j}$ depend on the dataset and model, they are not directly comparable across problems. To provide a meaningful scale, we normalize by the importance of a random noise feature, $\overline{\text{PFI}}_{\text{noise}}$, to provider a noise-normalized $\widetilde{\text{PFI}}_{j}$ in Eq. 11.
$\widetilde{\text{PFI}}_{j}=\frac{\overline{\text{PFI}}_{j}}{\overline{\text{PFI}}_{\text{noise}}}$ (11)
A $\widetilde{\text{PFI}}_{j}\approx 1$ implies that feature $j$ is no more predictive than random noise, while $\widetilde{\text{PFI}}_{j}>1$ indicates predictive power beyond noise. For example, $\widetilde{\text{PFI}}_{j}=5$ means that feature $j$ is five times as informative as the noise baseline. As shown in Figure 8, the anisotropy-based descriptors IAMS and QAMS dominate the model’s predictive performance, consistent with their direct link to polarization-dependent spectral asymmetries discussed in Sec. 2. The modest importance of geometric and chemical descriptors such as the normalized space group, density, and DAMS indicates that the model captures secondary contributions from the structural environment. The slight advantage of IAMS over QAMS likely reflects reduced uncertainty in quantities derived directly from computed spectra compared to oxidation-state–dependent charge estimates. e
3.0.3 Model Performance for Each Individual Metal
Figure 9: Per-element performance of the random forest regressor evaluated on the test set. Ordered from largest (top) to smallest (bottom) atomic number. (a) shows the coefficient of determination $R^{2}$ computed separately for each metal using the combined (unbalanced) training dataset, while (b) shows the corresponding $R^{2}$ values obtained using the balanced dataset. In both panels, the model is trained on all elements and evaluated on element-specific subsets of the test data.
To better understand the limitations of the model, we analyzed its performance on a per-element basis. Specifically, we computed the coefficient of determination, $R^{2}$, for each individual metal using subsets of the test data corresponding to a single element, while keeping the random forest regressor (RFR) trained on the full dataset. The resulting per-element $R^{2}$ values are shown in Figure. 9 (a).
For clarity, we explicitly define the coefficient of determination as
$R^{2}=1-\frac{\sum_{i}\left(y_{i}-\hat{y}_{i}\right)^{2}}{\sum_{i}\left(y_{i}-\bar{y}\right)^{2}},$ (12)
where $y_{i}$ are the true target values, $\hat{y}_{i}$ are the corresponding model predictions, and $\bar{y}$ is the mean of the target values for the subset under consideration.
Figure 10: Predicted versus true SAMS values for element-specific subsets of the test data. (a) shows the Sc subset, for which the model achieves a low coefficient of determination ($R^{2}=0.014$, $N=90$), indicating performance only marginally better than a mean predictor. (b) shows the Cu subset, where the model performs substantially better ($R^{2}=0.851$, $N=415$), with predictions closely following the ideal $y=x$ relationship (dashed line). In both panels, the model is trained on the full multi-element dataset and evaluated on element-specific test subsets.
Among all elements, Sc exhibits the poorest performance, with an $R^{2}$ value close to zero, indicating that the model provides little improvement over a naive mean predictor for this element. Figure 10 shows predicted versus true SAMS values for the Sc (a) and Cu (b) subsets of the test data. Showcasing the difference between a subset of the data that the model performs well on and one that it does not.
For Sc, the coefficient of determination is $R^{2}\approx 0.01$, which corresponds to a ratio of squared errors $\approx 0.99.$ This indicates that the model reduces the error relative to a mean predictor by only about 1%. The small but positive $R^{2}$ value suggests that the model predictions are slightly shifted relative to the Sc-specific mean, consistent with a bias toward the global mean $\overline{y}_{\mathrm{comb,true}}$ of the full multi-element dataset.
In contrast to Sc, Cu (b) exhibits excellent performance, with an $R^{2}$ of $0.851$, demonstrating that the model can accurately learn patterns when sufficient representative data is present.
Next, we investigate if the poor performance of Sc may be due to under representation of this element in the materials project dataset. To address this, we constructed a balanced data set with an equal number of compounds (444 compounds from each element, see Figure 6 (a)). The training and testing sets were also split such that there was an even distribution of each element, with 354 samples in the training, and 90 samples in the test set. The results of this are presented in Figure. 9 (b), where the performance for Sc significantly worsens, suggesting that its poor predictive behavior likely stems from intrinsic differences in the characteristics of Sc compared to the other 3dd transition metals rather than data imbalance.
One possible explanation for the poorer performance on Sc is that its chemical behavior differs significantly from the dataset’s majority. Unlike most 3dd transition metals, Sc exhibits nearly exclusive $3+$ oxidation, resulting in a $3d^{0}$ closed-shell configuration with weak ligand-field effects and minimal metal–ligand covalency. Consequently, its spectra are less anisotropic and structurally sensitive, differing fundamentally from the more covalent and partially filled $3d$ systems the model was trained on.
4 Conclusion
In this work we demonstrated that a random forest model is able to accurately predict a continuous anisotropy parameter using structural and chemical information extracted from crystallographic structures. This result provides an important step in creating new opportunities for machine learning driven materials discovery, as well as the design of materials with targeted spectroscopic properties. The model generalizes well to any crystalline 3dd transition metal structure (with the exception of Sc), but could easily be adapted to treat molecular systems as well by replacing or removing the normed space group feature.
It is also important to emphasize that the model should not be applied outside the scope in which it has been trained. While the approach demonstrates strong predictive capability within crystalline 3dd transition metal compounds, it is not applicable to materials with significantly different chemistry, dimensionality, or bonding environments (e.g., amorphous solids, $f$-electron oxides) without retraining or adaptation. In such cases, careful validation and retraining with an appropriately curated dataset are necessary before drawing conclusions.
Future work includes extending our machine learning model to predict anisotropy in other x-ray methods, such as XAS. In particular, many of the structural and chemical descriptors that drive anisotropy in XES are also expected to influence absorption edge features, suggesting that the framework developed here could be transferable across spectroscopies. It may be possible to apply a transfer learning approach so that the already trained XES anisotropy model can be fine-tuned for predicting anisotropy in a different technique, which could dramatically reduce the computational cost compared to training a new model from scratch.
Another important step is expanding the scope of the training data itself. Incorporating systems with controlled defects, surface terminations, or varying dimensionality would allow the model to capture a wider range of realistic environments, improving its applicability to experimental conditions. Beyond methodological development, coupling the model with high-throughput structure generation and screening pipelines could open new opportunities for accelerated materials discovery. For example, one could rapidly explore chemical substitutions, strain effects, or heterostructure geometries and identify candidates with large or tunable anisotropy.
5 Acknowledgments
JJK and JJR acknowledge support from the Theory Center for Materials and Energy Spectroscopies (TIMES) at SLAC funded by DOE BSE Contract DE-AC02-76SF00515. CAC was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2140004. GTS was supported by funding from the U.S. Department of Energy in the Nuclear Energy University Program under Contract No. DE-NE0009158. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the U.S. Department of Energy. This work was partly facilitated through the use of advanced computational infrastructure provided by the Hyak supercomputer system at the University of Washington.
6 Supplemental Information
6.1 Physical Motivation for Features
6.1.1 DAMS
The DAMS are useful for discriminating between crystalline structures that include null or anomalously small QAMS. In Figure 11 the local structure in (a) is clearly anisotropic and should therefore produce a large SAMS signal. However, from (b) we can see that there is a situation for which the the QAMS can made to equal zero (for a given distortion of the ligands), specifically $\alpha=1/\sqrt{2}$.
This example highlights a flaw in the QAMS that is inherent to its construction, namely that it’s possible to produce a zero QAMS for a clearly anisotropic system. Here we use a relatively simple example, but the same result can be found for other geometries, and the condition will also depend on the normalization constant $m$. However, this failing can be offset by the inclusion of the DAMS, which can only be made to be zero for $\alpha=0$, and therefore provide orthogonal information to the model that can inform it local anisotropy that it would not be able to infer from the QAMS alone.
Figure 11: (a) Distorted square planar structure with a TM ion (silver) lying at (0, 0, 0) and four ligands lying along the $x$, $y$ axis that have all been translated positively by $\alpha$ along the $z$ axis. The charge $q_{n}$ of each ligand has been assigned as 1 for convienience. (b, c) The calculation of the (QAMS, DAMS) for an arbitrary normalization constant $m$.
6.2 Permutation Feature Importance
Permutation Feature Importance (PFI) begins with the definition of a performance metric. In our case we use the negative mean squared error (NMSE):
$\text{NMSE}(y,\hat{y})=-\frac{1}{n}\sum_{i=1}^{n}\big(y_{i}-\hat{y}_{i}\big)^{2},$
where $y_{i}$ are the true values, $\hat{y}_{i}$ are the predicted values, and $n$ is the number of samples. We adopt NMSE following the convention in scikit-learn, where higher scores indicate better performance. Since MSE is minimized, its negation ensures that values closer to zero represent better performance, while more negative values indicate worse performance [sklearn_permutation_importance].
Let $\text{NMSE}_{0}$ denote the baseline score of the model on the unpermuted dataset. For each feature $j$, we construct a permuted dataset $X^{\pi(j)}$ by randomly shuffling the entries in column $j$ while leaving all other columns unchanged. We then recompute the score
$\text{NMSE}_{j}=\text{NMSE}(y,\hat{y}^{\pi(j)}).$
The PFI for feature $j$ in a single permutation trial is defined as
$\text{PFI}_{j}=\text{NMSE}_{0}-\text{NMSE}_{j}.$
This process is repeated $R$ times with different random permutations, and the average importance is taken as
$\overline{\text{PFI}}_{j}=\frac{1}{R}\sum_{r=1}^{R}\Big(\text{NMSE}_{0}-\text{NMSE}_{j}^{(r)}\Big).$